OpenStudy (anonymous):

Rohan draws a card from a deck of 52 cards. Each time when he draws a card, he notes the result and places the card back in the deck. He does this experiment 100 times. The table given below shows the occurrence of each type of card. What is the difference between the probabilities of drawing a black card and a red card?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

Probablity of a black card = 60/100 = 3/5 or 0.6 Probability of a red card = 40/100 = 2/5 or 0.4

OpenStudy (anonymous):

@mathsmind would u plse tell me if it is correct or not

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (shubhamsrg):

Seems fine.

OpenStudy (anonymous):

now i have to subrtact

OpenStudy (shubhamsrg):

yep.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thank u @shubhamsrg